OpenStudy (anonymous):

how to solve & explain 48/X=12

OpenStudy (anonymous):

are you looking for x? 48/12 = x

OpenStudy (anonymous):

multiply both sides by x to get 48=12x, then divide both sides by 12 to get 48/12=x

OpenStudy (anonymous):

10 amunptah Medals 0 (48/X)=12= 48/x * x/1=x*12= 48=12x=48/12=12x/12= X=4 thanks ludwig 457